The mean and variance of a binomial distribution are 4 and .(4/3)Find P(X>=1)

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
From the given data np=4....(1),npq=4/3.........(2)
dividing 2 by 1 we get q=3/4 and p=1-q=1/4 and n=16
P(x\tiny \geq 1)=1-P(x=0)=1-(3/4)16=1-0.01=0.99
Thanks and Regards
Shaik Aasif
askIITians faculty